OpenStudy (anonymous):

If sin A , cos A , tan A are in Geometric Progression . then cos^3A+cos^2A = ??

OpenStudy (amistre64):

well what does a geometric progression tell us about how the terms are formed?

OpenStudy (anonymous):

cos a=rsina tan a=r^2sina

OpenStudy (amistre64):

correct, another way to look at it may be from the center out in this case: sin, sin cot, sin cot cot (= tan)

OpenStudy (amistre64):

had some competing thoughts lol

OpenStudy (amistre64):

cos tan = sin cos = cos cos sin/cos^2 = tan

OpenStudy (amistre64):

sin/cos^2 needs to equal cot giving us cos^2 = sin tan

OpenStudy (amistre64):

.... need paper lol

OpenStudy (amistre64):

cos^3+cos^2 cos^2(cos + 1) (1-sin^2)(cos + 1) (1 + sin) (1 - sin) (cos + 1) just some reworkings to try to get it into some useable format

OpenStudy (amistre64):

is there any more context to the question? any idea as to what its looking for as a solution? is it a proof, or something else?

OpenStudy (amistre64):

do we need to solve for A?

OpenStudy (anonymous):

answer is given as 1. thats about it. we need to find cos^2A+cos^3A if sin,cos, tan are in gp

OpenStudy (amistre64):

yeah, solving for A is paramount then

OpenStudy (amistre64):

sinA * r = cosA if r = cot(A) sinA*r^2 = tanA if r^2 = secA or a geometric mean might help too cosA = sqrt(sinA tanA) cos^2A = sin^2A/cosA cos^3A = sin^2A cos^3A = 1 - cos^2A got it

OpenStudy (anonymous):

ohohoh. yeah. didnt see that. i was trying to break down cos^2 A simultaneously. xD Thankyouu. :)

OpenStudy (amistre64):

youre welcome .... the geometric mean made this much simpler